OpenStudy (zenmo):

Very small question (will probably take a second), help me understand this textbook example. ~Calculus Limits & Derivatives.

OpenStudy (zenmo):

How is that 0?

OpenStudy (thomas5267):

Why is \(\lim_{x\to a}(x-a)\neq 0\)? \(a-a=0\).

OpenStudy (zenmo):

So, substitute 0 in \[\lim_{x \rightarrow a}(x-a) \] to get a-a=0?

OpenStudy (zenmo):

substitute 0 for a*

OpenStudy (thomas5267):

Since \(f(x)=x-a\) is continuous you can do that.

OpenStudy (zenmo):

Ok. Thanks! :)

OpenStudy (thomas5267):

Continuity of \(f(x)=x-a\) can be proven using \(\epsilon\text{-}\delta\) definition of limit.